Author Archives: Isaac Bakerman

140801 UCSC Observations

Isaac arrived at ISB at 9:50am. Upon arrival the doors to the collector had been deployed and the big fan as well as the little fan were on. A sample of approximately 50ml had been collected in the sample jar. The sample was then removed and stored in a labeled bag. The collector was then disassembled and sprayed lightly with 409 followed by a thorough rinse with DI water. The disassembled collector parts were additionally sprayed with 409 followed by a thorough rinse with DI water. The collector was then reassembled and a blank was taken. The  system was then turned back on and reset for the next fog event.

140730 LML Observations

Isaac arrived at LML at 9:25 am. The doors to the fog collector were closed due to the previous intentional shut down of the system. Several tests were ran on the system to further elucidate the determining factor in the system malfunction. The first test ran observed strictly the humidity and temperature sensor setting on the RPi. Upon increase of humidity on the sensor the system still ran and was fully operational. The second test involved running the fog collection program on the RPi. In this case when the humidity was increased the solenoids did not fire. However, the big fan did for a short while and then turned off. At this point the program had crashed. The system was then turned off and then a third test was administered. this test was preformed by executing the test function on the RPi. The test revealed that neither solenoid fired but the big fan was working and then  shut off. Upon completion of observations the system was shut off.

140728 LML Observations

Isaac arrived at LML at 9:30am both doors on the fog collector had not opened and no sample had been collected. Upon closer examination the system had appeared to have been turned off prior to arrival. It was later determined that this was done during the previous day when the collector was serviced. The collector was turned back on and immediately the system began to run the program to collect the fog. the system at this point was operational however when the humidity measurement was increased by blowing on the sensor the system began to initiate the process to collect fog. Sounds for the solenoids to fire were noted but the solenoids did not fire and the system froze. The system was turned off and a test was run. The test showed that the solenoids again did not fire but the big fan did turn on for awhile until it automatically went off. A manual test was then administered to test the firing of the solenoids and the big fan. The results of this test showed that the solenoids did not fire but the big fan did fire. The system was then turned off in order to determine further causes of the crashing.

140724 UCSC Observations

Isaac arrived at the top of ISB at 9:30 am. The doors to the collector had opened but no sample had been collected. The collector was disassembled and the body was sprayed with 409 followed by a thorough rinse with DI water. The collector components were additionally sprayed with 409 followed by a thorough DI water rinse. The collector was then reassembled for the next fog event.

140720 LML Observations

Isaac arrived at the collector at 9:50 am. The collector doors had not opened and no sample had been collected. Upon arrival the small fan was running. The system was then tested to ensure the program was operational. The test showed that the system was working as intended. The collector was then disassembled and its components were rinsed with DI water. The collector was then rinsed with DI water and the collector was then reassembled and reset for the next fog event.

140719 LML Observations

Isaac arrived at LML at 9:30 am. The doors to the fog collector had been deployed although there was no collected sample. The collector components were then removed and cleaned with strictly DI water. The body was then sprayed with a very small amount of 409 followed by a rinse with DI water. The collector was then reassembled and reset for the next fog event.

140718 UCSC Observations

Isaac arrived at the UCSC location at 9:50am. No sample was visible in the collection jar although the doors to the fog collector had opened. The collector components were then removed and sprayed initially with 409 followed by thorough rinsing with DI water. The collector was then sprayed with 409 and DI water as well. The collector jar was then washed with strictly DI water. The collector was then reassembled and a blank was taken. The collector was then reset for the next fog event.

140714 LML Observations

Isaac arrived at the LML site at 10:00am. No sample had been collected from the previous evening’s fog event. It was noted though that there was a high volume of insects on the strand frames and in the drip tube assembly. The collection jar was removed and cleaned with strictly DI water. The collector was then disassembled and its components were sprayed initially with 409 followed by a thorough rinse with DI water. The collector was additionally sprayed with 409 and DI water. The collector was then reassembled and reset for the next fog event.

140711 UCSC Observations

Isaac arrived at ISB at 10:30am.  A sample had been collected from the fog event the previous evening. This sample jar was removed and later stored in the UCSC cold room. The collector was then disassembled and the components were cleaned with 409 followed by DI water. The collector was then reassembled and a blank was taken. The system was then reset for the next fog event.

140709 UCSC Observations

Isaac arrived at the fog collector on top of ISB at 9:30am. The doors to the collector had opened and a sample was collected in the sample jar measuring approx 200ml. The jar was removed and stored in the PSB cold room upon completion of cleaning. The body of the collector, the string frames, and drip tube assembly were sprayed lightly with 409 followed by a thorough rinse with DI water. A blank was then taken and a new sample jar put in place. The system was reset for the next fog event.